### v1

# SlowFast-SCI: Slow-Fast Deep Unfolding Learning for Spectral Compressive Imaging

Abstract
Humans learn in two complementary ways: a slow, cumulative process that builds broad, general knowledge, and a fast, on-the-fly process that captures specific experiences. Existing deep-unfolding methods for spectral compressive imaging (SCI) mirror only the slow component—relying on heavy pre-training with many unfolding stages—yet they lack the rapid adaptation needed to handle new optical configurations. As a result, they falter on out-of-distribution cameras, especially in bespoke spectral setups unseen during training. This depth also incurs heavy computation and slow inference. To bridge this gap, we introduce SlowFast-SCI, a dual-speed framework seamlessly integrated into any deep unfolding network beyond SCI systems. During slow learning, we pre-train or reuse a priors-based backbone and distill it via imaging guidance into a compact fast-unfolding model. In the fast learning stage, lightweight adaptation modules are embedded within each block and trained self-supervised at test time via a dual-domain loss—without retraining the backbone. To the best of our knowledge, SlowFast-SCI is the first test‐time adaptation–driven deep unfolding framework for efficient, self‐adaptive spectral reconstruction. Its dual‐stage design unites offline robustness with on‐the‐fly per‐sample calibration—yielding over 70% reduction in parameters and FLOPs, up to 5.79 dB PSNR improvement on out‐of‐distribution data, preserved cross‐domain adaptability, and a 4× faster adaptation speed. In addition, its modularity integrates with any deep-unfolding network, paving the way for self-adaptive, field-deployable imaging and expanded computational imaging modalities. Code and models are available at https://github.com/XuanLu11/SlowFast-SCI
1 Introduction
Figure 1: (a) Previous SCI reconstruction method vs. our SlowFast-SCI. (b) The PSNR-FLOPs-Params analysis comparing the proposed SlowFast-SCI with latest state-of-the-art methods.
Humans are endowed with a complementary learning system mcclelland1995there ; hong2024slowfast that operates on two distinct timescales. A slow process in the neocortex gradually integrates diverse experiences into a coherent world model, enabling robust prediction of action outcomes schwesinger55review . In parallel, a fast process in the hippocampus rapidly encodes episodic details, supporting agile adaptation to novel circumstances tulving1983elements . Integrating such a dual-speed mechanism into neural networks enables rapid adaptation, i.e., swift recalibration to new domains or unseen inputs—thereby enhancing both flexibility and generalization.
This dual-speed learning paradigm inspires advances in spectral compressive imaging (SCI), a computational photography technique that aims to reconstruct rich hyperspectral information from compressed measurements. The predominant SCI architecture, coded aperture snapshot spectral imaging (CASSI) gehm2007single_cassi ; wagadarikar2008single_cassi_sparsity ; meng2020endtsa_cassi , employs a spatially coded mask and a dispersive element to modulate, shift, and integrates spectral bands onto a two-dimensional sensor zhang2022herosnet , achieving high-throughput, cost-effective, and bandwidth-efficient acquisition. The resulting hyperspectral images (HSIs) deliver rich spatial–spectral signatures that significantly enhance scene characterization beyond conventional RGB imagery. Consequently, SCI has found broad applications across computer vision and remote sensing, e.g., object detection van2010tracking_object ; kim20123d_object , medical imaging lu2014medical ; Meng_2020_OL_SMEM_medical , and land-cover classification borengasser2007hyperspectral_remote ; yuan2017hyperspectral_remote .
Based on the CASSI acquisition model, a broad spectrum of reconstruction algorithms has been proposed to recover a three-dimensional hyperspectral data cube from its two-dimensional measurement. Traditional model-based methods employ handcrafted priors—such as sparsity wagadarikar2008single_cassi_sparsity ; kittle2010multiframe_sparsity , total variation rudin1992nonlinear_tv ; wang2015dual_tv , and low-rank constraints liu2018rank_lowrank ; zhang2019computational_lowrank —but they require careful parameter tuning and often yield suboptimal quality and slow convergence. The advent of deep learning introduced end-to-end networks that significantly improve both reconstruction accuracy and speed mst ; cst ; dauhst ; hdnet ; meng2020gap ; meng2020endtsa_cassi ; bisci ; Dong_2023_CVPR_rdluf ; Zhang_2024_CVPR_ssr ; Zhang_2024_CVPR_dpu .
In particular, deep unfolding frameworks (DUNs) unroll iterative optimization into trainable layers, thereby integrating model-driven regularization with data-driven learning meng2020gap ; huang2021deep ; Zhang_2024_CVPR_dpu ; dauhst ; Dong_2023_CVPR_rdluf . However, these methods depend on extensive pre-training over large, paired datasets—an impractical requirement given the scarcity and acquisition cost of real HSI data, reflecting only the slow component of learning. Moreover, to achieve state-of-the-art performance, deep unfolding approaches often employ a large number of sequential unfolding blocks; each block must process the full high-dimensional hyperspectral volume, which greatly increases computational and memory burdens. This level of complexity not only complicates training and inference but also makes such architectures impractical for real-world applications and deployment on resource-constrained or edge devices.
Per-sample adaptation—analogous to human fast learning—is essential for robust SCI reconstruction under scarce HSI data. While recent advances such as self-supervised fine-tuning and test-time adaptation enable rapid calibration, applying them directly to deep unfolding networks proves problematic: full-network fine-tuning destabilizes sequential stages and linear probes cannot accommodate layer-wise inference. SAH-SCI liu2024sahsci supports general self-supervised tuning but lacks runtime adaptation, and recent test-time adaptation (TTA) methods zhao2024adapnet ; Qin_2023_CVPR ; quan2022dual have yet to address hyperspectral SCI.
To overcome these challenges, we introduce SlowFast-SCI, the first test-time adaptation-driven deep unfolding framework for efficient, self-adaptive spectral reconstruction. In the slow learning stage, we pre-train a deep unfolding backbone that fuses model-based priors with learnable parameters to acquire robust, general reconstruction capabilities. In the fast learning stage, lightweight Fast TTA modules (FAST-TTAMs) are inserted into each unfolding block and self-supervisedly fine-tuned at test time, instantly aligning the model to new measurements while preserving the frozen backbone. By uniting offline robustness with on-the-fly adaptation, SlowFast-SCI achieves superior cross-domain generalization without additional labels.
Furthermore, DUNs typically rely on numerous sequential stages to ensure high reconstruction fidelity, but this depth comes at the cost of heavy computation and slow inference. To address this, we apply an imaging mechanism-guided model distillation step that compresses the pre‐trained backbone into a streamlined “fast unfolding” variant with significantly reduced complexity. We then integrate the same lightweight adaptation modules into this distilled network and optimize only these modules during inference using a dual‐domain self‐supervised loss quan2022dual . With the backbone frozen, this design delivers swift test‐time tuning and efficient inference, mirroring the human ability to rapidly leverage distilled knowledge for new experiences. We summarize our contributions as follows:
  • We propose a unified dual-speed learning framework, SlowFast-SCI, which augments any deep unfolding network with offline pre-training for robust reconstruction and lightweight on-the-fly adaptation to new spectral configurations. To the best of our knowledge, this is the first test-time adaptation–driven, self-adaptive spectral unfolding framework.
  • We develop a lightweight self-supervised adaptation module embedded in each unfolding stage to enable rapid per-sample calibration at inference.
  • A model distillation pipeline that compresses the pre-trained backbone into a fast unfolding variant, reducing parameter count and FLOPs by over 70% while preserving adaptability.
  • Evaluation on standard Harvard, ICVL and real CASSI measurements, showing up to 5.79 dB PSNR gain on out‐of‐distribution setups and a $4\times$ speedup in adaptation time.
2 Related Work
Hyperspectral Image Reconstruction. Related work in hyperspectral image reconstruction comprises two main streams. Traditional optimization‐based methods employ hand‐crafted priors—such as sparsity wagadarikar2008single_cassi_sparsity ; kittle2010multiframe_sparsity , total variation rudin1992nonlinear_tv ; wang2015dual_tv , and low‐rank constraints liu2018rank_lowrank ; zhang2019computational_lowrank —to iteratively solve the inverse problem; while interpretable, they are limited by prior expressiveness and suffer from slow convergence. More recently, deep learning and hybrid approaches have accelerated SCI reconstruction mst ; cst ; dauhst ; hdnet ; Dong_2023_CVPR_rdluf ; Zhang_2024_CVPR_dpu , including end‐to‐end networks that map 2D measurements directly to hyperspectral volumes, plug‐and‐play frameworks that insert learned denoisers into iterative solvers, and deep unfolding architectures—e.g., DAUHST dauhst , RDLUF Dong_2023_CVPR_rdluf , and DPU Zhang_2024_CVPR_dpu —that unroll optimization iterations into trainable layers. These learning‐based methods deliver state‐of‐the‐art accuracy and speed but depend heavily on large paired datasets of hyperspectral images and measurements, making them brittle under optical variations or spectral distribution shifts at test time. In contrast, our work introduces a complementary paradigm: a self‐supervised, test‐time adaptation framework that enables rapid per‐sample calibration without additional labeled data, thereby improving robustness and generalization across unseen scenes.
Test-Time Adaptation. TTA aims to adapt a pre-trained model during inference to better align it with test data, improving performance without access to labeled test sample sun2020test_tta ; darestani2022test_tta . In image reconstruction, TTA methods have been introduced to fine-tune pre-trained models at test time quan2022dual ; Qin_2023_CVPR ; zhao2024adapnet . AdaptNet zhao2024adapnet embeds self-supervised fine-tuning blocks into deep unfolding networks, yielding strong performance in CS-MRI and SV-CT, but its considerable depth leads to slow adaptation. In contrast to previous inefficient approaches, we propose a lightweight adaptation module and a concise distillation pipeline, which together enable efficient test-time adaptation while exploiting the strengths of adapter architectures.
Model Distillation. Model distillation buciluǎ2006model_distill ; hinton2015distilling is a widely adopted technique for compressing large models into smaller ones by transferring knowledge from a pre-trained teacher to a student network. While early methods focused on soft-label supervision for classification tasks, subsequent work has extended distillation to various forms, including intermediate feature matching romero2014fitnets , attention transfer Zagoruyko2017AT , and data-free distillation zhu2021data . Distillation has also been applied to self-supervised learning caron2021emerging , multi-task learning li2020knowledge , and test-time adaptation wang2021tent . Existing distillation techniques primarily target semantic or classification tasks, leaving inverse problems—especially SCI‐based image reconstruction—largely unexplored. We address this by introducing a structural‐knowledge distillation framework that transfers imaging priors from a pre-trained backbone into a lightweight, fast‐unfolding student model. As a result, the student inherits the teacher’s robustness while remaining both efficient and adaptable during test‐time inference.
Figure 2: Illustration of SlowFast-SCI. The left side depicts the slow learning process, pre-training on the source data with a deep unfolding network. The right side shows the fast learning process, where FAST-TTAM rapidly adapts the fast unfolding network to align the model with target scenes at inference time. The far right illustrates the details of FAST-TTAM, which takes both spatial and spectral information into account during fast learning.
3 Method
SlowFast-SCI tackles two complementary challenges in SCI reconstruction with dual-speed learning. In the slow learning stage (Sec. 3.1), we overcome the ill-posed inverse problem and scarce HSI training data by pre-training a physics-guided unfolding backbone that fuses model priors with learnable parameters to learn robust, general reconstruction capabilities. In the fast learning stage (Sec. 3.2), we address domain shifts and per-sample variability by inserting lightweight spectral adapters into each unfolding block and self-supervisedly fine-tuning only these adapters at test time—instantly calibrating to unseen measurements while keeping the backbone frozen. Furthermore, we propose an imaging-guided distillation framework that condenses the hybrid architecture into a compact “fast unfolding” network, augmented with tailored spectral adaptation blocks. This design enables efficient, label-free reconstruction across domains, mirroring the rapid application of distilled expertise in human learning.
3.1 Slow Learning
Figure 3: The schematic of CASSI system.
Mathematical Model of CASSI. SCI is a snapshot modality that captures a full hyperspectral data cube in a single coded measurement by jointly modulating and integrating across all spectral bands. As illustrated in Fig. 3, the CASSI implements SCI as follows. Let a hyperspectral patch consist of $N_{\lambda}$ bands $\{X_{i}\}_{i=1}^{N_{\lambda}}$, each $X_{i}\in\mathbb{R}^{H\times W}$. Each band is first modulated by a known binary mask $M$, then the modulated frames are shifted along the detector by d(i1)d(i-1) pixels and summed to yield the compressed measurement $Y(u,v)=\sum_{i=1}^{N_{\lambda}}M\odot X_{i}\bigl(u,\,v+d(i-1)\bigr)+N(u,v),$ where $\odot$ denotes the Hadamard product, $N$ is additive sensor noise, and $(u,v)$ index spatial coordinates. By vectorizing the 3D hyperspectral cube into $x\in\mathbb{R}^{nN_{\lambda}}$ and the 2D measurement into $y\in\mathbb{R}^{n}$ with $n=H\bigl(W+d\,(N_{\lambda}-1)\bigr),$ the acquisition model can be written in matrix–vector form as
$\mathbf{y}=\Phi\mathrm{\mathbf{x}}+\mathbf{n},$ (1)
where $\Phi\in\mathbb{R}^{n\times nN_{\lambda}}$ encodes the combined masking and shifting operations. SCI reconstruction then consists of recovering the hyperspectral vector $\mathbf{x}$ from the single compressed measurement $\mathbf{y}$ given the known sensing matrix $\Phi$.
Deep Unfolding Network-driven Slow Learning. Traditional optimization-based spectral reconstruction guarantees convergence and interpretability but tends to be brittle; CNNs and transformers add flexibility but require large labeled datasets—unrealistic in spectral imaging. Deep unfolding unrolls a MAP optimizer into a trainable network, marrying solver stability with learned representational power to achieve high-fidelity reconstructions from limited data. Therefore, we adopt this MAP-based unfolding as our slow-learning backbone for rapid adaptation.
In general, the deep unfolding networks formulate SCI reconstruction as a Bayesian problem, solving Eq. (1) under a unified MAP framework:
$\hat{\mathrm{\mathbf{x}}}_{\mathit{map}}=\arg\min_{\mathrm{\mathbf{x}}}\log P(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})=\arg\min_{\mathrm{\mathbf{x}}}\log P(\mathrm{\mathbf{y}}|\mathrm{\mathbf{x}})+\log P(\mathrm{\mathbf{x}}).$ (2)
In particular, the original signal could be estimated by minimizing the following energy function:
$\hat{\mathrm{\mathbf{x}}}=\arg\min_{\mathrm{\mathbf{x}}}\frac{1}{2}\|\mathrm{\mathbf{y}-\Phi\mathrm{\mathbf{x}}}\|_{2}^{2}+\lambda R(\mathrm{\mathbf{x}}),$ (3)
where $\frac{1}{2}\|\mathrm{\mathbf{y}-\Phi\mathrm{\mathbf{x}}}\|^{2}$ and $R(\mathrm{\mathbf{x}})$ represent the data fidelity term and the image prior term respectively and $\lambda$ is a trade-off hyperparameter. Among many solutions for Eq. (3), the HQS method can formulate the problem as an iterative scheme as follows:
$\mathrm{\mathbf{r}}_{k+1}$ $=\mathrm{\mathbf{x}}_{k}+\mathbf{\Phi}^{\mathsf{T}}\left[(\mathrm{\mathbf{y}}-\mathbf{\Phi}\mathrm{\mathbf{x}}_{k})/(\mu+\mathbf{\Phi}\mathbf{\Phi}^{\mathsf{T}})\right],$ (4)
$\mathrm{\mathbf{x}}_{k+1}$ $=\arg\min_{\mathrm{\mathbf{x}}}\frac{\mu}{2}\|\mathrm{\mathbf{x}}-\mathrm{\mathbf{r}}_{k+1}\|_{2}^{2}+\lambda R(\mathrm{\mathbf{r}}),$ (5)
where $\mu$ is a penalty parameter. Based on the iterative scheme above, one can have a deep unfolding network with many stages. Within each stage, the $\mathbf{x}$-sub problem can be regraded as a denoising problem, that is, to predict the original HSI signals $\mathrm{\mathbf{x}}$ from their noisy counterparts $\mathrm{\mathbf{r}}=\mathrm{\mathbf{x}}+\epsilon$, which is solved by a learnable denoising network.
3.2 Fast Learning
Our slow learning phase (Sec. 3.1) endows the deep unfolding network with robust physics‐informed priors and learnable parameters. Yet when faced with a new spectral domain, naively fine‐tuning every unfolding block demands as much data and compute as the original training. To overcome this, we observed that each denoiser—when linearized—collapses into an optimal Wiener filter qin2023ground ; zhao2024adapnet , which led us to insert lightweight adaptation blocks into each stage and self‐supervisedly fine‐tune only these per test sample (Sec. 3.2.1). Then, we distill the pre-trained network into a nimble ‘fast unfolding’ variant and finally introduce a customized self-supervised imaging guidance loss on the adapters (Sec. 3.2.2), delivering rapid inference and seamless cross-domain adaptation.
3.2.1 Fast Learning via Wiener-Filter–Inspired Adapter Modules
Let $\mathcal{D}(\cdot,\omega)$ denote a denoising network with parameters $\omega$. Then the Eq. (5) is transformed into the following form:
$\mathrm{\mathbf{x}}_{k+1}=\mathcal{D}(\mathrm{\mathbf{r}}_{k+1},\omega_{k+1}).$ (6)
Training $\mathcal{D}(\cdot,\omega)$ to minimize prediction error means minimizing the mean squared error (MSE) by updating the parameters $\omega$:
$\min_{\theta}\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x}),\mathbf{\epsilon}\sim p(\mathbf{\epsilon})}\{\|\mathcal{D}(\mathrm{\mathbf{r}},\omega)-\mathbf{x}\|_{F}^{2}\}\approx\min_{\theta}\sum_{i}\|\mathcal{D}(\mathrm{\mathbf{r}}_{(i)},\omega)-\mathbf{x}_{(i)}\|_{F}^{2}\}.$ (7)
How to understand the statistical properties of general CNNs as a form of Bayesian inference is an open question. For the sake of discussion, we consider a CNN consisting of only linear convolution layers, which does not contain nonlinear activation functions and bias terms, and its expression is:
$\mathcal{D}(\mathrm{\mathbf{r}},\omega)=\mathcal{H}\otimes\mathrm{\mathbf{r}},$ (8)
where $\otimes$ represents the convolution operator and $\mathcal{H}$ is the filter of $N_{\lambda}$ channels. Let $\mathcal{H}_{w}$ denote the optimal filter, Wiener filter, we have, for each channel $\lambda_{i}$,
$\widehat{\mathcal{H}}_{w}[k_{1},k_{2},\lambda_{i}]=\frac{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}[k_{1},k_{2},\lambda_{i}]|^{2}\}}{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}[k_{1},k_{2},\lambda_{i}]|^{2}\}+\mathbb{E}_{\mathbf{\epsilon}}\{|\widehat{\mathbf{\epsilon}}[k_{1},k_{2},\lambda_{i}]|^{2}\}},$ (9)
where $k_{1},k_{2}]$ denotes the frequency of the image in the frequency domain after discrete Fourier transformation (DFT). The derivation details of Eq. (9) are given in the supplementary material. Similarly, for a specific test sample $\tilde{\mathrm{\mathbf{r}}}=\tilde{\mathrm{\mathbf{x}}}+\tilde{\mathrm{\mathbf{\epsilon}}}$, its individual optimal wiener filter for each channel $\lambda_{i}$ is given by
[…]
Pre-training a linear CNN on source data effectively learns the Wiener estimator for that distribution—derived from the image and noise second‐order statistics in the frequency domain (Eq. (9)). Since the Wiener solution for a specific target scene (Eq. (10)) only differs modestly from the source estimator, a compact, supplementary convolutional filter can bridge this gap, aligning the pre-trained source filter with the target‐scene Wiener filter.
3.2.2 Self-Adaptation-Driven Fast-Learning Unfolding
Imaging-Guided Fast Unfolding Distillation. The process of model distillation is shown in Fig. 2. During model distillation, the pre-trained model, with its parameters frozen, serves as the teacher model, while the two-stage model with the same architecture as the pre-trained model is used as the student model. The dataset used for distillation is a dataset that the pre-trained model has not seen during supervised slow learning. Firstly, mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times(W+d\times(N_{\lambda}-1)}$ are entered into the teacher model to obtain the results $\{\mathrm{\mathbf{X}}_{th,i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$, which will be used as a label to supervise the training of the student model. Then mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ are input into the student model to obtain the results $\{\mathrm{\mathbf{X}}_{st,i}\}_{i=1}^{N_{d}}$. We update the student model by minimizing the following distillation loss:
$\mathscr{L}_{dis}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\|\mathrm{\mathbf{X}}_{st,i}-\mathrm{\mathbf{X}}_{th,i}\|_{F}^{2}.$ (11)
The inference speed of the two-stage student model obtained by model distillation is much faster than the original pre-trained model, and its performance on test data is comparable to that of the original pre-trained model.
Adapter-Enhanced Unfolding Network Architecture. Motivated by the discussion in Sec. 3.2.1, we propose a two-stage unfolding network where each stage is attached by a single lightweight module, namely FAST-TTAM, which takes both spatial and spectral information into account. The distilled unfolding network, is a typical unfolding network for SCI reconstruction. For each stage of the distilled unfolding network, we concatenate the adjustable lightweight operator to the end of the frozen stage. The FAST-TTAM is a lightweight operator with one update step, one parameter $\mu$ estimation network, and two convolution blocks, the spatial convolution block and the spectral convolution block. The structure of parameter $\mu$ estimation network is shown in the upper right of Fig. 2, which is used to estimate the parameter $\mu$. The spatial convolution block consists of two convolution layers with $3\times 3$ filter and one $\mathrm{ReLU}$ function, while the spectral convolution block only includes one convolution layer with $1\times 1$ filter.
Self-Supervised Training and Test-Time Adaptation. Consider the distilled unfolding network with adaptive layers $\mathcal{F}_{DA}(\cdot;\omega^{*},\theta)$:
$\mathcal{F}_{DA}(\cdot;\omega^{*},\theta):\mathrm{\mathbf{y}},\mathbf{\Phi}\rightarrow{\mathrm{\mathbf{x}}},$ (12)
where $\omega^{*}$ denotes the parameters of the distilled model frozen in self-supervised training and test-time adaptation and $\theta$ denotes the parameters of the adapt layers to be updated. Given the same dataset $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ as in model distillation, we update $\theta$ by minimizing the following self-supervised loss:
$\mathscr{L}_{sst}=\mathscr{L}_{m}+\omega_{1}\mathscr{L}_{ei}+\omega_{2}\mathscr{L}_{iu}+\omega_{3}\mathscr{L}_{tv}.$ (13)
[…]
$\mathbb{R}^{N}=\mathrm{im}(\mathbf{\Phi}^{H})\oplus\mathrm{ker}(\mathbf{\Phi}),$ (14)
where $\mathrm{im}(\mathbf{\Phi}^{H})=\{\mathbf{\Phi}^{H}\mathrm{\mathbf{y}},\mathrm{\mathbf{y}}\in\mathbb{R}^{M}\}$ and $\mathrm{ker}(\mathbf{\Phi})=\{\mathrm{\mathbf{x}}\in\mathbb{R}^{N},\mathbf{\Phi}\mathrm{\mathbf{x}}=\mathbf{0}\}$. To solve the severe linear inverse problem of SCI during test-time adaptation, the key is how to adapt the model for prediction with small prediction errors in both $\mathrm{im}(\mathbf{\Phi}^{H})$ and $\mathrm{ker}(\mathbf{\Phi})$ efficiently at the test time. To address this, we propose a test-time self-supervised loss function $\mathscr{L}_{tta}$:
$\mathscr{L}_{tta}:=\mathscr{L}_{im}+\lambda\mathscr{L}_{ker},$ (15)
with a pre-defined hyperparameter $\lambda$. $\mathscr{L}_{im}$ and $\mathscr{L}_{ker}$ concern the prediction errors in $\mathrm{im}(\mathbf{\Phi}^{H})$ and $\mathrm{ker}(\mathbf{\Phi})$ respectively. For the prediction in $\mathrm{im}(\mathbf{\Phi}^{H})$, we use the measurement $\mathrm{\mathbf{y}}$ to train our network.
$\mathbf{\Phi}\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta)\rightarrow\mathrm{\mathbf{y}}\approx\mathbf{\Phi}\mathrm{\mathbf{x}}+\mathbf{n}.$ (16)
For the prediction in $\mathrm{ker}(\mathbf{\Phi})$, we train our model using the output $\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta)$ as label and $\mathbf{\Phi}T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta))$ as input, where $T(\cdot)$ represents the data augmentation (shifts, rotations, etc.).
$\mathcal{F}_{DA}(\mathbf{\Phi}T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta);\omega^{*},\theta))\rightarrow T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta);\omega^{*},\theta)).$ (17)
Table 1: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI (SlowFast-SCI1 and SlowFast-SCI2 represent the pre-trained model we used were DAUHST and DPU, respectively.) and SOTA methods on 10 simulation scenes of Harvard dataset.
Method Reference Params (M) FLOPs (G) S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 Avg
TSA-Net meng2020endtsa_cassi ECCV 2020 42.20 91.58
30.54
0.795
25.76
0.557
30.27
0.750
22.24
0.422
36.71
0.845
27.62
0.820
27.51
0.610
24.83
0.644
27.23
0.570
22.79
0.490
27.55
0.650
MST mst CVPR 2022 1.92 25.60
32.67
0.857
30.28
0.699
31.20
0.798
26.21
0.515
38.63
0.938
30.40
0.837
32.80
0.751
29.34
0.791
32.40
0.732
27.85
0.670
31.18
0.759
CST cst ECCV 2022 2.86 36.77
33.17
0.882
32.29
0.789
32.90
0.845
27.74
0.595
39.44
0.960
29.29
0.871
33.60
0.812
31.53
0.867
34.00
0.811
28.60
0.733
32.26
0.817
BIRNAT bisci TPMAI 2023 4.19 1973.66
35.58
0.914
35.23
0.884
34.93
0.864
30.23
0.721
42.38
0.976
35.97
0.917
36.28
0.896
33.62
0.907
36.87
0.888
30.48
0.807
35.16
0.877
RDLUF Dong_2023_CVPR_rdluf CVPR 2023 1.81 115.34
37.06
0.929
34.36
0.863
36.49
0.891
30.04
0.687
44.21
0.983
36.13
0.929
36.78
0.885
33.27
0.903
37.04
0.885
32.02
0.826
35.74
0.878
SSR-L Zhang_2024_CVPR_ssr CVPR 2024 5.18 78.93
36.25
0.921
33.60
0.838
35.38
0.880
29.01
0.666
43.23
0.978
36.41
0.924
35.79
0.863
33.57
0.890
35.87
0.860
30.03
0.745
34.91
0.857
DAUHST-SAH liu2024sahsci ECCV 2024 3.73 49.79
35.10
0.912
33.09
0.812
34.84
0.870
28.18
0.593
42.82
0.975
34.30
0.918
34.60
0.850
32.09
0.867
34.54
0.823
29.86
0.736
33.94
0.836
DPU-SAH liu2024sahsci ECCV 2024 3.12 54.44
36.93
0.932
34.41
0.853
36.04
0.884
30.19
0.697
43.43
0.980
35.51
0.930
36.20
0.867
33.61
0.895
36.80
0.874
31.02
0.779
35.41
0.869
DAUHST-9stg
(slow learning)
dauhst
NeurIPS 2022 6.15 79.50
34.95
0.914
31.14
0.758
34.67
0.873
26.34
0.532
42.56
0.978
33.62
0.920
32.64
0.798
30.14
0.845
32.51
0.770
27.81
0.694
32.64
0.808
SlowFast-SCI1 Ours 1.41 21.84
35.52
0.917
36.23
0.868
33.91
0.850
31.06
0.694
42.25
0.971
34.18
0.910
37.89
0.871
36.05
0.909
37.80
0.863
33.16
0.813
35.81
0.867
DPU-9stg
(slow learning)
Zhang_2024_CVPR_dpu
CVPR 2024 2.85 49.26
36.90
0.932
34.36
0.852
36.02
0.884
30.14
0.695
43.41
0.980
35.45
0.930
36.14
0.866
33.56
0.894
36.72
0.873
30.97
0.777
35.37
0.868
SlowFast-SCI2 Ours 0.70 15.17
37.15
0.930
41.93
0.949
36.15
0.879
37.94
0.865
44.09
0.978
37.65
0.934
45.88
0.971
42.25
0.965
45.52
0.960
43.09
0.962
41.16
0.939
Figure 4: Simulated HSI reconstruction comparisons (middle) on Harvard. The left shows the scene measurement and the right shows the spectral curves corresponding to the selected region. Please zoom in for better view.
4 Experiments
4.1 Experiment Setup
Datasets. The pre-trained model used in slow learning was trained on the CAVE park2007multispectralcave dataset. We adopted two datasets Harvard chakrabarti2011statistics_harvard and ICVL arad2016sparse_icvl with randomly clipped patches of size $256\times 256\times 28$ for simulation experiments. Harvard/ICVL dataset contains 67/180 images for training and 10/20 for testing. For real datasets, we used 5 scenes with spatial size $660\times 660$ and 28 bands captured in TSA-Net meng2020endtsa_cassi , then cropped data blocks of size $256\times 256\times 28$ for testing.
Comparison Methods. We compare our SlowFast-SCI on synthetic data and real scenes with state-of-the-art (SOTA) reconstruction methods including TSA-Net meng2020endtsa_cassi , MSTmst , CSTcst , BIRNATbisci , RDLUFDong_2023_CVPR_rdluf , SSR-L Zhang_2024_CVPR_ssr , SAH-SCI liu2024sahsci , DAUHST dauhst , and DPU Zhang_2024_CVPR_dpu .
Model Distillation and FAST-TTAM Training. In model distillation, we distilled two typical deep unfolding networks with 9 stages, DAUHST darestani2022test_tta and DPU Zhang_2024_CVPR_dpu . The distillation epoch and batch size were set to 200 and 2, respectively. We used the Adam 2014Adam optimizer for both networks. We set the initial learning rate as $4\times{10}^{-4}$ and adopted the Cosine Annealing learning rate scheme loshchilov2016sgdr to implement training. We trained the FAST-TTAM for 100 epochs. The batch size was set to 2 and the learning rate was fixed at $1\times{10}^{-4}$. The $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$ were set to 0.7, 0.4 and 0.001 in Eq. (13).
Test-Time Adaptation. The test-time adaptation was conducted over 50 iterations for Harvard and 10 iterations for ICVL, with a constant learning rate of $1\times{10}^{-3}$ and $\lambda=0.7$ throughout all iterations.
4.2 Quantitative Results
Comparison on Harvard and ICVL.
Tab. 1 and 2 summarize PSNR, SSIM, parameter count, and FLOPs for SlowFast-SCI1 (slow-learning backbone: DAUHST) and SlowFast-SCI2 (slow-learning: DPU) against SOTA methods. SlowFast-SCI2 achieves 41.16dB/0.939 on Harvard and 39.79dB/0.963 on ICVL, outperforming all competitors. SlowFast-SCI1 yields 35.81dB/0.867 and 37.68dB/0.950, improving its backbone by 9.71%/7.30% on Harvard and 15.12%/3.26% on ICVL. Both variants use <25% of the parameters and <33% of the FLOPs of their backbones. Compared to SAH-SCI, SlowFast-SCI2 delivers 16.24%/8.06% and 6.79%/1.16% gains on Harvard and ICVL, respectively, with lower resource costs; SlowFast-SCI1 shows similarly strong improvements.
Table 2: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI and SOTA methods on 10 simulation scenes of ICVL dataset.
Method Reference Params (M) FLOPs (G) S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 Avg
TSA-Net meng2020endtsa_cassi ECCV 2020 42.20 91.58
26.22
0.720
28.50
0.736
28.02
0.777
26.32
0.819
31.10
0.736
28.23
0.750
20.91
0.783
25.14
0.762
28.74
0.700
27.25
0.817
27.05
0.760
MST mst CVPR 2022 1.92 25.60
27.07
0.877
31.89
0.909
31.80
0.883
30.43
0.900
35.79
0.931
32.22
0.893
24.05
0.829
27.00
0.842
34.93
0.907
26.30
0.815
30.15
0.879
CST cst ECCV 2022 2.86 36.77
26.61
0.871
31.42
0.904
31.97
0.885
29.36
0.881
36.14
0.935
32.00
0.884
23.26
0.792
26.35
0.833
35.24
0.910
26.72
0.846
29.91
0.874
BIRNAT bisci TPMAI 2023 4.19 1973.66
34.78
0.934
38.41
0.955
36.93
0.951
38.52
0.957
42.17
0.970
37.40
0.945
30.35
0.926
32.59
0.906
37.73
0.932
34.87
0.938
36.37
0.941
RDLUF Dong_2023_CVPR_rdluf CVPR 2023 1.81 115.34
36.67
0.953
41.16
0.970
40.79
0.975
40.93
0.971
45.72
0.983
39.34
0.959
34.12
0.956
34.73
0.928
39.87
0.951
37.27
0.960
39.06
0.961
SSR-L Zhang_2024_CVPR_ssr CVPR 2024 5.18 78.93
35.77
0.938
41.14
0.968
39.39
0.966
41.20
0.970
45.53
0.982
39.69
0.960
33.24
0.947
34.72
0.922
38.90
0.942
37.30
0.946
38.69
0.954
DAUHST-SAH liu2024sahsci ECCV 2024 3.73 49.79
29.94
0.902
34.24
0.937
34.50
0.936
32.88
0.934
39.74
0.962
34.98
0.929
26.37
0.893
29.82
0.879
37.23
0.926
28.39
0.878
32.81
0.918
DPU-SAH liu2024sahsci ECCV 2024 3.12 54.44
33.90
0.940
39.55
0.966
38.12
0.971
39.82
0.967
45.26
0.981
39.36
0.960
28.31
0.919
32.63
0.918
39.76
0.951
35.85
0.950
37.26
0.952
DAUHST-9stg
(slow learning)
dauhst
NeurIPS 2022 6.15 79.50
29.49
0.902
34.26
0.941
34.43
0.938
32.83
0.938
39.94
0.966
35.07
0.932
25.59
0.890
29.61
0.883
37.45
0.932
28.59
0.883
32.73
0.920
SlowFast-SCI1 Ours 1.41 21.84
35.48
0.936
39.99
0.964
38.91
0.963
39.73
0.964
43.86
0.975
37.55
0.946
32.23
0.938
34.31
0.921
38.40
0.934
36.35
0.952
37.68
0.950
DPU-9stg
(slow learning)
Zhang_2024_CVPR_dpu
CVPR 2024 2.85 49.26
33.88
0.939
39.51
0.966
38.09
0.971
39.76
0.966
45.21
0.981
39.32
0.960
28.29
0.919
32.61
0.918
39.74
0.951
35.81
0.950
37.22
0.952
SlowFast-SCI2 Ours 0.70 15.17
38.59
0.962
41.33
0.970
41.54
0.979
42.15
0.972
45.49
0.981
39.22
0.957
35.72
0.957
36.16
0.937
39.45
0.949
38.26
0.962
39.79
0.963
Figure 5: Simulated HSI reconstruction comparisons (middle) on ICVL. The left shows the scene measurement and the right shows the spectral curves corresponding to the selected region. Please zoom in for better view.
4.3 Qualitative Results
Simulation Data Results. Figs. 4 and 5 depict the simulation HSI reconstruction comparisons between our method and other SOTA methods on Harvard and ICVL datasets, respectively. It can be seen that the other models produce poor reconstruction results with a lot of noise and artifacts, since they lack fast learning on similar data. The results of our SlowFast-SCI in different bands show that the proposed dual speed framework can correct the problem of weak generalization ability of SCI reconstruction due to insufficient training data. This is because our method enhances the generalization of the model to extract more spectral information through self-supervised fast learning and can adapt the model to align it with target scenes at inference. Besides, the spectral density curves show that the reconstruction results of SlowFast-SCI are more similar and correlated with the GT, which proves our method is capable of improving the spectral-dimension consistency of the model.
Real-Scene HSI Reconstruction. We evaluate SlowFast-SCI on real-world HSI using CAVE and KAIST choi2017highkaist , pre-training a two-stage DPU backbone with real CASSI masks under the protocols of Zhang et al. Zhang_2024_CVPR_dpu , Meng et al. meng2020endtsa_cassi , and Cai et al. dauhst . To simulate realistic noise, we inject 11-bit shot noise during both slow- and fast-learning. At test time, raw measurements serve as self-supervised labels for adaptation. As shown in Fig. 6, SlowFast-SCI recovers finer details and suppresses noise more effectively than SOTAs, demonstrating strong real-world potential.
We also provide additional visual comparisons of SlowFast-SCI against SOTAs on scenes from the Harvard, ICVL datasets and real dataset—see the Supplementary Material for details.
Figure 6: Real-scene HSI reconstruction results. Shown are reconstructions of two real-world scenes across 2 selected spectral channels (out of 28), using various state-of-the-art methods. Compared to others, SlowFast-SCI restores finer details and better suppresses noise, highlighting its effectiveness in real-data scenarios.
4.4 Ablation Studies
Fast Unfolding-Stage Configuration Analysis. We assess the effect of different unfolding stages on performance using the Harvard dataset (Tab. 3), where FAST-TTAM denotes the unfolding layer is attached by the adaptation module. Using the 9-stage DPU as a baseline (case a), results from cases (a, c, f) show that both 1-stage and 2-stage variants preserve the key characteristics of the original model. However, cases (d, f) and (e, g) show that the 1-stage variant learns less effectively than the 2-stage network. Thus, despite its faster inference, the 1-stage model is discarded in favor of the better-performing 2-stage variant that enables a balance between reconstruction quality and inference speed.
Table 3: Ablation on the stages of fast unfolding network, FAST-TTAM, and Test-time adaptation.
Case Stages FAST-TTAM TTA PSNR SSIM Params (M) FLOPs (G)
(a) 9 $\times$ $\times$ 35.37 0.868 2.85 49.26
(b) 9 $\times$ 34.93 0.850 3.16 67.92
(c) 1 $\times$ $\times$ 35.67 0.876 0.32 5.39
(d) 1 $\times$ 38.39 0.912 0.35 7.63
(e) 1 38.39 0.903 0.35 7.63
(f) 2 $\times$ $\times$ 35.75 0.876 0.64 10.86
(g) 2 $\times$ 39.97 0.932 0.70 15.17
(h) 2 41.16 0.939 0.70 15.17
Figure 7: Ablation on iteration times at inference.
Fast Learning Ablation. We analyze the performance of FAST-TTAM and TTA in the fast learning. Tab. 3 (c, d) and (f, g) show that the proposed FAST-TTAM can effectively transfer the distilled model to the target domain through self-supervised training. Besides, case (g, h) in Tab. 3 shows self-supervised test-time adaptation can further align the fast unfolding variant with target scenes.
Iteration Times at Inference. This experiment shows the relationship between iteration times and performance gains through TTA. Fig. 7 shows that as the number of iterations increases, the PSNR improves, and the required inference time also increases accordingly. Only a few iterations are needed to achieve impressive performance, which proves the effectiveness of our method.
5 Conclusion
We have introduced SlowFast-SCI, to our knowledge the first test-time adaptation–driven, self-adaptive spectral unfolding framework. By adopting a dual-speed paradigm—distilling a priors-driven backbone into a compact “fast unfolding” network during offline training, then embedding lightweight, self-supervised adapters for on-the-fly calibration—our method achieves over 70% parameter and FLOPs reduction, up to 5.79 dB PSNR improvement on out-of-distribution data, and a 4× faster adaptation runtime, all while preserving robust cross-domain performance.
Crucially, SlowFast-SCI’s modularity allows it to be seamlessly integrated into any deep unfolding architecture, unlocking real-time hyperspectral reconstruction on resource-constrained platforms—ranging from aerial drones and point-of-care scanners to distributed environmental sensors. We believe this work paves the way for a new class of self-adaptive, field-deployable imaging systems and look forward to exploring extensions to other computational imaging modalities and richer adaptation objectives.
Limitation and Future Work: SlowFast-SCI’s per-sample adapters, despite their small footprint, add inference overhead that may hinder use on ultra–low-power or hard real-time platforms. Our study is limited to disperser-based CASSI; applying and validating the dual-speed approach on other SCI architectures and unfolding models remains crucial. Future directions include hardware-aware adapter designs, early-exit adaptation schemes, and extending the framework to snapshot multispectral, fluorescence, tomographic imaging and general image restoration.